### v2

# Homogenized C. elegans Neural Activity and Connectivity Data

Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans). This is particularly interesting as this model system provides a path to bridge the gap between structure and function, from nervous system connectivity to physiology. However, the many existing physiology datasets, both recording and stimulation, as well as connectome datasets, are in distinct formats, requiring extra processing steps before modeling or other analysis can commence. Here we present a homogenized dataset of neural activity, including during stimulation, compiled from 11 neuroimaging experiments and from 10 connectome reconstructions. The physiology datasets, collected under varying experimental protocols, all measure neural activity via calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling frequency. The resulting dataset includes neural recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy (EM) reconstructions, all contain the entire nervous system of the worm, preprocessed into a graph of connections across the neurons. Using our collection of datasets is facilitated through easy data sharing on HuggingFace. We believe that our joint dataset of physiology and connectivity will facilitate modeling, for example in terms of recurrent neural network or transformer architectures, making it easier to check how well models generalize across animals and labs.
1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience and machine learning. While neural dynamics in the brain can often be complex and difficult to measure in larger organisms, smaller systems like C. elegans offer a unique opportunity to create comprehensive models that bridge functional activity with known connectomic structure. C. elegans has a relatively small nervous system of 300 neurons [17], and both its neural dynamics and complete synaptic connectome have been extensively studied and mapped out [24, 3, 25]. This makes it an ideal organism for studying how neural dynamics arise from connectivity and how this relationship could inform both neuroscience and AI models of neural function.
With its transparent body and well-characterized neural architecture, C. elegans provides an excellent model for whole-brain functional imaging using calcium fluorescence sensors [26]. Calcium indicators act as proxies for local computation, allowing researchers to capture neural activity at single-neuron resolution across many neurons simultaneously [6]. Additionally, since the full connectome of C. elegans has been mapped via electron microscopy [24, 3, 25], combining these two datasets — neural dynamics and connectivity structure — allows for a deep investigation of how network connectivity influences neural function. Unification of such datasets can also contribute to the broader goal of understanding how neural activity propagates through networks, ultimately informing both the development of AI-based neural systems and providing insights into how these mechanisms generalize to more complex organisms [8].
However, the existing datasets from calcium imaging studies of C. elegans are often recorded under various experimental conditions, with different subsets of neurons labeled and differing sampling rates. This makes it challenging to perform cross-study comparisons or build unified models.
To address these challenges, we have created two datasets, the first combining neural dynamics from 11 calcium maging studies and the second combining neural connectivity from 3 connectome studies of C. elegans [24, 3, 25]. This dataset integrates both structure and function, making it an invaluable resource for computational modelers interested in building neural network models grounded in real biological data. The primary goal of this dataset is to facilitate the training of computational models of C. elegans neural dynamics, which can be used to study how neural function is related to structure. The dataset includes preprocessed neural activity data that has been normalized, resampled to a common time step, and smoothed using a causal exponential kernel. It also includes graph-based representations of the connectome, with detailed information about neuron positions and synaptic connections, facilitating the development of models that combine structure and function. We are releasing these datasets as open-source resources on the HuggingFace platform to ensure broad accessibility to researchers in both neuroscience and machine learning. By unifying these datasets and making them accessible, we hope to encourage the development of computational models that can capture the structure-function relationships in small nervous systems.
2 Methods
2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions and files) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data in a standardized format.
The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor. The former manages processing of neural activity data as measured by calcium fluorescence imaging [23]. The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy [24].
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The former handles the extraction and processing of the raw calcium data from each of our 11 neural activity source datasets into standardized neural activity data that are saved as compressed files. The latter processes the raw connectome data from connectome source files into compressed graph tensors that are saved as compressed in a data format amenable to downstream processing by graph neural network packages. There are 10 distinct connectome source files as shown in Table 3 but they are all derived from 4 primary source publications [24, 3, 25, 14].
Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 26, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as showb in Table 1). All animals were hermaphrodites at developmental stages from L4 to adulthood.
Table 1: Calcium Fluorescence Neural Activity Datasets Metadata. Metadata for calcium fluorescence neural activity datasets collected from various sources. Each dataset includes information about the experimental conditions, the number of worms, the average number of labeled and recorded neurons, and the range of labeled neurons observed.
Source Dataset Name Database Link Data Files Num. Worms Mean Num. Neurons (labeled, recorded) Num. labeled neurons (min, max)
Kato2015 [9] osf.io/2395t/ WT_Stim.mat 12 (42, 127) (31, 51)
Nichols2017 [12] osf.io/kbf38/ let.mat 44 (34, 108) (23, 43)
Skora2018 [15] osf.io/za3gt/ WT_.mat 12 (46, 129) (39, 55)
Kaplan2020 [9] osf.io/9nfhz/ Neuron2019_Data_ *.mat 19 (36, 114) (23, 51)
Yemini2021 [26] zenodo.org/records/ 3906530 _Activity_OH.mat 49 (110, 125) (33, 179)
Uzel2022 [18] osf.io/3vkxn Uzel_WT.mat 6 (50, 138) (46, 58)
Dag2023 [13] tinyurl.com/github Dag2023 data/swf702_with_ id/*.h5 7 (100, 143) (87, 110)
Atanas2023 [5] wormwideweb.org YYYY-MM-DD-.json/h5 42 (88, 136) (64, 115)
Leifer2023 [14] tinyurl.com/drive Leifer2023 exported_data.tar.gz 108 (64, 68) (19, 98)
Lin2023 [11] tinyurl.comdropbox Lin2023 run_prfrd_data.mat 577 (8, 8) (1, 22)
Venkatachalam2024 [22] chemosensory-data.worm.world 2022_herm_.zip 22 (187, 187) (185, 189)
2.2.1 Neural Data Processing
The entire pipeline is implemented in the NeuralBasePreprocessor class, which handles dataset-specific preprocessing tasks, including file format loading, trace extraction, and metadata creation.
Table 2: Notation Used in the Neural Data Processing Pipeline. Description of the mathematical notation used throughout the neural data preprocessing pipeline. Variables represent worm-specific indices, neural activity data structures, temporal properties, neuron counts, and preprocessing parameters.
Symbol Description
$k$ Worm index variable within a single dataset
$\mathbf{X}^{(k)}$ Neural activity data for a single worm $k$
$T_{k}$ Number of time points of data recorded for a worm $k$
$D$ Number of neurons in C. elegans
$\mathbf{\mu},\mathbf{\sigma}$ Mean and standard deviation over time of neural activity data for each neuron
$\alpha$ Optional smoothing hyperparameter
$\mathbf{M}^{(k)}$ Binary mask generated for each worm using labeled neurons applied to neural activity data
The neural activity data is preprocessed using a custom pipeline (Figure 1) which follows these steps:
1. Data Download and Extraction. The raw data from each dataset is downloaded from its respective source and extracted. The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking.
We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C. elegans. The neural activity at any time point $t$ is denoted as $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, which represents the activity snapshot across the $D$ neurons.
3. Neuron Trace Normalization. The first preprocessing step is to normalize or z-score the raw (i.e. from the source datasets) neural activity across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
where $\mathbf{\mu},\mathbf{\sigma}\in\mathbb{R}^{D}$ are the mean and standard deviation over time for each neuron.
4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}_{\text{norm}}[t]+(1-\alpha)\mathbf{X}^{(k)}_{\text{norm}}[t-1]$
where $\alpha$ is a smoothing hyperparameter set in preprocess.yaml. For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed. Calcium fluorescence data inherently reflects low frequency neural dynamics. Smoothing, which acts further low pass over the raw data, is not needed.
5. Resampling. We resample the data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
$t^{\prime}=\left\lfloor\frac{t\cdot\Delta t_{\text{original}}}{\Delta t}\right\rfloor$
The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed as:
$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}[t]+\left(\mathbf{X}^{(k)}[t+1]-\mathbf{X}^{(k)}[t]\right)\cdot(t^{\prime}-t)$
This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics.
Importantly, the resampling step is the last preprocessing step necessary to create the neural activity data matrix for each worm:
$\mathbf{X}^{(k)}\leftarrow\mathbf{X}_{\text{resampled}}^{(k)}$
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled and measured for worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not measured.
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$ enabling in its potential use in downstream analyses that require filtering out unlabeled neurons. Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
Figure 1: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm. The mask does not modify the data but is retained as metadata to aid future analyses (e.g., selecting only labeled neurons).
Algorithm 1 Neural Data Preprocessing Pipeline
[…]
for each dataset $\mathcal{D}_{\mathrm{raw,source}}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
     for each worm $k$ in $\mathcal{D}_{i}$ do
         Extract calcium fluorescence traces from source files: $\mathbf{X}^{(k)}_{\text{raw}}$
         Normalize traces: $\mathbf{X}^{(k)}_{\text{norm}}\leftarrow\text{normalize}(\mathbf{X}^{(k)}_{\text{raw}})$
         (Optional) Smooth traces: $\mathbf{X}^{(k)}_{\text{smooth}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{smooth}},\Delta t)$
         Processed neural data: $\mathbf{X}^{(k)}\leftarrow\mathbf{X}^{(k)}_{\text{resample}}$
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\mathbf{X}^{(k)}_{j}\mid\mathbf{X}^{(k)}_{j}\neq\emptyset,\,j\in[1,D]\}$
         Store processed neural data and labeled neuron mask: $\mathbf{X}^{(k)}$ and $\mathbf{M}^{(k)}$
     end for
     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
end for
Combine the processed datasets
2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets that capture the full neural wiring diagram, including both chemical synapses and gap junctions. Table 5 summarizes these datasets, which vary in annotation precision and recording techniques.
Table 3: Open-Source Connectome Datasets Metadata.Metadata for open-source C. elegans connectome datasets. Each dataset was standardized into the graph_tensor format, listing the corresponding graph tensor file and the total number of edges (synaptic and electrical connections) per dataset.
Source Dataset Name Graph Tensor File Number of Edges
Cook2019 [3] graph_tensors_cook2019.pt 46837
White1986 (whole) [24] graph_tensors_white1986_whole.pt 46603
White1986 (n2u) [24] graph_tensors_white1986_n2u.pt 45808
White1986 (jsh) [24] graph_tensors_white1986_jsh.pt 45759
White1986 (jse) [24] graph_tensors_white1986_jse.pt 44901
Witvliet2020 (7) [25] graph_tensors_witvliet2020_7.pt 46057
Witvliet2020 (8) [25] graph_tensors_witvliet2020_8.pt 46107
OpenWorm2023 [21] graph_tensors_openworm.pt 47683
Randi2023 [14] graph_tensors_funconn.pt 50521
Chklovskii2011 [19] graph_tensors_chklovskii.pt 46655
The connectome data is represented as a directed graph where neurons are nodes and edges represent synaptic or gap junction connections.
Figure 2: Graphical Representation of the Connectome Data. Graph-based structure used for the C. elegans connectome data. Neurons are represented as nodes, with connections (edges) classified as either chemical synapses or gap junctions. The neuron features, such as positional coordinates, neuron class, and type, are stored as node attributes, while edge attributes capture synaptic weights and connection types. This graph-based structure enables connectivity-informed modeling of neural activity. The differently colored nodes arbitrarily represent the presence of different types of neurons within the C. elegans nervous system (e.g. motor, inter & sensory neurons). The tables correspond to the bolded edge and node. The attr(e_1) feature for a single edge is the tuple of chemical synapse and gap junction weights.
2.3.1 Graph Tensor Data Processing
The data processing is handled by the ConnectomeBasePreprocessor class, which standardizes the neuron indices, processes edge attributes, and constructs the final graph tensors.
Table 4: Attributes in Graph Tensor Format. Attributes and their descriptions in the standardized graph tensor format. Includes neuron identities, edge connectivity types, synaptic weights, and spatial coordinates.
Symbol Description
$k$ Number of labeled neurons in a dataset
$l$ Number of labeled edges in a dataset
$N_{\text{pre}}^{k}$ Enumerated neuron identities for presynaptic neurons
$N_{\text{post}}^{k}$ Enumerated neuron identities for postsynaptic neurons
$T^{k}$ Connection type binary values (gap junction vs. chemical synapse)
$E$ Presynaptic neuron, postsynaptic neuron pairs which form edges
$G^{l}$ Gap junction weights for $l$ edges
$C^{l}$ Chemical synapse weights for $l$ edges
$x,y,z$ Cartesian coordinate $x,y,z$ position values for each neuron
1. Column Data Format.
Within each source connectome, presynaptic and postsynaptic neurons were specified along with chemical synapse and gap junction weight values. In the majority of the source connectome files, the format tabular for $k$ neurons with a column for the presynaptic neuron identity denoted by $N_{\text{pre}}^{k}$, a column for postsynaptic neuron identity denoted by $N_{\text{post}}^{k}$, connection type (gap junction vs. chemical synapse) denoted by $T^{k}$ with binary value elements (0 or 1), and the number of synapses, $N_{\text{weight}}^{k}$.
Table 5: Column Data Format for Connectome Dataset. Example of tabular data structure used in connectome datasets. Columns describe the origin and target neurons, connection type (gap junction or chemical synapse), number of connections, and associated neurotransmitters. The example shown is the first and last few rows of the OpenWorm [21] connectome source file
. Origin Target Type Number of Connections Neurotransmitter ADAL ADAR GapJunction 1 Generic_GJ ADAL ADFL GapJunction 1 Generic_GJ ADAL AIBL Send 1 Glutamate ADAL AIBR Send 2 Glutamate ADAL ASHL GapJunction 1 Generic_GJ VC1 VC2 Send 3 Serotonin_Acetylcholine VC1 VC2 GapJunction 6 Generic_GJ VC1 VC3 Send 1 Serotonin_Acetylcholine VC1 VC3 GapJunction 2 Generic_GJ VC1 VD1 GapJunction 1 Generic_GJ
2. Matrix Connectome Data Format.
The exception to the tabular connectome source dataset format with distinct columns for presynaptic and postsynaptic neuron identities was the functional connectome from [14] and the connectome from [3]. These connectomes were in $N\times N$ matrix format where $N$ is the number of neurons. Neurons were arranged alphabetically. The $(i,j)$ element within the matrix represents the synaptic strength proportional to elicited response of surrounding stimulated neurons for the functional connectome from [14]. Similarly, the $(i,j)$ element within the matrix represents the number of synapses between neurons $i$ and $j$ for the connectome from [3]
3. Extraction.
Each connectome has a custom class created which inherits from the parent ConnectomeBasePreprocessor class. This was due to each connectome source being formatted differently from each other requiring custom loading and preprocessing from different file types in different formats.
Within each custom subclass, we implemented a preprocess function which was designed to open the source dataset respective file type and extract the relevant data to construct the final graph tensor object. Each preprocess function ends with the graph tensor object being saved in tensor file format via the save_graph_tensor() function call. Depending on the structure of the source dataset files, each custom preprocess function contained code responsible for opening and iterating through the different blocks of the file to extract relevant connectivity and edge weights data.
We used the torch_geometric library to instantiate a Data graph object which standardized the data structure used by the various connectome sources.
4. Graph Tensor Formatting
For each source connectome dataset, we have a list of neurons $N$ and $l$ edges $E\in\{N_{pre},N_{post}\}^{l}$. The list of edges between neurons which form the connectome are saved as the edge_index attribute of the graph tensor object. Additionally, synaptic weights for both gap junctions and chemical synapses were extracted from each source connectome dataset and saved as the edge_attr attribute of the graph tensor object. The gap junction weights, denoted by $G\in R^{l}$, and chemial synapse weights, denoted by $C\in R^{l}$, were appended together to produce the edge_attr tensor attributes. These weightages provide a standardized denotion of neuron influence to better construct graph conditioned, informed, and constrained neural activity models.
5. Cross-Dataset Pairing
The source connectome datasets did not contain all of the necessary information to produce the graph tensor object which was subsequently saved as a tensor file. Specifically, information such as the neuron type used by certain chemical synapse connections between neurons were not provided with every single source connectome dataset. As a result, we obtained connectivity information from each source dataset and paired that with neuron information associated with each chemical synapse edge. Additionally, other attributes were also shared across datasets which were not present in the original source datasets such as $x,y,z$ position of neurons. This cross-dataset pairing strategy allows for a more comprehensive and standardized graph tensor object format across different source datasets.
2.3.2 Consensus Connectome Analysis
Analysis across the different connectome source datasets was performed to extract commonalities between specific source connectome chemical synapse and gap junction connections. Specifically, we created a consensus connectome that, for every neuron-to-neuron chemical synapse or gap junction connection across all source datasets, selects the maximum weight value.
3 Discussion
In this paper we present a normalized and standardized dataset for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code blocks we have provided are composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
This unified neural and connectome dataset provides a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. However, while this dataset offers numerous advantages, there are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. One notable clarification is that while C. elegans is often said to have 302 neurons, we refer to it as having 300 neurons in this work because two of these cells, CANL and CANR, are now classified as “end organs” rather than neurons. This shift in classification is supported by [17], which redefined the roles of these cells based on regulatory gene expression, leaving 300 functional neurons to form the core of the nervous system.
One of the primary limitations arises from the use of calcium fluorescence imaging as the method for measuring neural activity. Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6]. Calcium signals lag behind the actual electrical activity due to the slower dynamics of calcium ion concentration changes compared to the fast voltage changes associated with action potentials. This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost. Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that is critical for certain neural circuits [23, 20]. This highlights the need to complement calcium imaging with other techniques, such as voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity.
Additionally, while the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and potentially across the lifespan of a single individual. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [3, 14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity. This limitation means that the relationship between structure and function is inferred across different animals, potentially missing individual-specific relationships between connectivity and function, despite phenotypic matching. The generation of the C. elegans connectome, as described by [24], provides a comprehensive structural map of synaptic connections, though understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons due to neuromodulator & neuropeptide release which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [14].
We opted not to apply smoothing to the neural activity traces for two principal reasons. First, it is likely that the original data creators have already undertaken a degree of preprocessing on their raw microscopy or imaging data, which often includes smoothing techniques to enhance signal quality. Introducing additional smoothing could risk over-processing the data, thereby obscuring critical neural dynamics. Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself. This filtering effectively diminishes high-frequency components, which are vital for capturing rapid neural events. By refraining from further smoothing, we aim to preserve the integrity and richness of the original data, allowing for a more nuanced understanding of the biological phenomena under investigation.
Resampling was chosen as a strategy to make the data comparable across experiments with different temporal resolutions. We resampled the neural activity traces to a common time step of approximately $\Delta t\approx 0.333$ seconds, which was chosen as a balance between temporal resolution and minimizing data loss due to downsampling from experiments with higher frame rates. However, this decision introduces its own limitations. There was a tradeoff between attenuating high-frequency information from datasets with higher sampling rates versus incorrectly hallucinating data from datasets with lower sampling frequencies. For experiments that originally had much higher temporal resolution, downsampling risks losing potentially valuable high-frequency dynamics. Conversely, for lower-resolution datasets, resampling can artificially inflate the temporal resolution without actually adding more meaningful information [11]. This was a necessary compromise to allow for cross-study comparisons but should be considered when interpreting the results.
On the computational side, while this dataset serves as a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using this dataset for machine learning models. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the neural dynamics data, affected by the limitations of calcium imaging, might bias models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional datasets (e.g., voltage imaging) to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. Graph neural networks (GNNs), for instance, are well-suited to handle the structural connectivity data, while recurrent neural networks (RNNs) and other time-series models can be applied to the neural activity data to capture temporal dependencies [9]. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the dataset’s open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
In conclusion, while this dataset offers significant opportunities for advancing our understanding of neural systems, its limitations should be carefully considered in both biological and computational contexts. Calcium imaging, structural connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used. As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C. elegans and the more complex neural architectures found in larger organisms [13, 12].
4 Conclusion
We present a standardization protocol implemented in a code repository for the preprocessing and integration of 11 C. elegans calcium fluorescence datasets and 3 primary connectome sources [24, 3, 25]. The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data. Each dataset has a class implementation that inherits from a respective parent class, which has been adapted to accommodate varying raw data formats. Once loaded, the calcium fluorescence data is further masked, normalized, resampled, and smoothed.
This unified dataset offers an accessible and standardized resource, streamlining access to data recorded under diverse experimental conditions and formats. The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data. Additionally, the full code repository containing the preprocessing code can be accessed on GitHub at worm-graph.
Such standardization facilitates the development of models for the entire C. elegans nervous system, allowing researchers to study the interplay between structure, via the connectome, and function, via calcium fluorescence data. Through this integration, we hope to provide further insights into how neural function emerges from structural connectivity and how these dynamics might generalize to more complex nervous systems.
The open-source nature of this resource aims to further contribute to both neuroscience and AI research, accelerating efforts to model not only C. elegans but also organisms with larger and more complex nervous systems. By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural dynamics modeling and the understanding of nervous systems across species.