### v3

# Homogenized C. elegans Neural Activity and Connectivity Data

Abstract
There is renewed interest in modeling and understanding the nervous system of the nematode Caenorhabditis elegans (C. elegans), as this small model system provides a path to bridge the gap between nervous system structure (connectivity) and function (physiology). However, existing physiology datasets, whether involving passive recording or stimulation, are in distinct formats, and connectome datasets require preprocessing before analysis can commence. Here we compile and homogenize datasets of neural activity and connectivity. Our neural activity dataset is derived from 11 C. elegans neuroimaging experiments, while our connectivity dataset is compiled from 9 connectome annotations based on 3 primary electron microscopy studies and 1 signal propagation study. Physiology datasets, collected under varying protocols, measure calcium fluorescence in labeled subsets of the worm’s 300 neurons. Our preprocessing pipeline standardizes these datasets by consistently ordering labeled neurons and resampling traces to a common sampling rate, yielding recordings from approximately 900 worms and 250 uniquely labeled neurons. The connectome datasets, collected from electron microscopy reconstructions, represent the entire nervous system as a graph of connections. Our collection is accessible on HuggingFace, facilitating analysis of the structure-function relationship in biology using modern neural network architectures and enabling cross-lab and cross-animal comparisons.
1 Introduction
Understanding neural dynamics and their relationship to underlying structure remains one of the key challenges in neuroscience. Recent advances in computational modeling and data acquisition have reinvigorated interest in the neural system of the nematode Caenorhabditis elegans (C. elegans), a model organism uniquely suited for bridging the gap between neural connectivity and function. With its fully mapped synaptic connectome [35, 36], consisting of approximately 300 neurons [3], and its extensively studied neuronal dynamics [23, 37], C. elegans provides an ideal platform for investigating how neural structure constrains or influences physiology.
Recent research has focused on leveraging the C. elegans connectome for generative models of neural development [30] and linking whole-brain activity to synaptic connectivity [4]. These studies exemplify its potential for elucidating how neural signaling is determined or constrained by connectivity and for inspiring biologically grounded computational frameworks [26]. Furthermore, the transparent body of C. elegans, combined with its well-characterized neural architecture, makes it an excellent candidate for whole-brain functional imaging using calcium fluorescence sensors [37]. Calcium imaging enables the capture of neural dynamics at single-neuron resolution across many neurons simultaneously [6], serving as a proxy for neuronal activity. Although calcium fluctuations may primarily reflect local processes in neurites rather than whole-neuron computations [24], compact neurons like those in C. elegans likely rely on such localized signaling mechanisms, making this an informative approach.
The combination of calcium imaging datasets with the full connectome, mapped via electron microscopy [35, 3, 36], offers an unprecedented opportunity to investigate how connectivity influences neural function. Unified datasets of structure and function can provide valuable insights into neural activity propagation, informing both the development of artificial intelligence (AI)-based neural simulations and the understanding of mechanisms that generalize to more complex organisms [8].
However, calcium imaging datasets from C. elegans experiments vary widely in their experimental conditions, labeled neurons, and sampling rates, complicating cross-study comparisons and unified modeling efforts. Similarly, connectome datasets—whether physical or functional—require significant preprocessing to facilitate analysis or computational modeling.
To address these challenges, we have created two integrated datasets. The first aggregates neural dynamics from 11 calcium imaging studies, and the second compiles connectome data from three electron microscopy wiring studies [35, 3, 36] and one functional signal propagation study [14]. Together, these datasets combine structure and function, offering an invaluable resource for computational modelers aiming to construct biologically grounded neural network models.
Our preprocessing pipeline standardizes neural activity data through normalization, resampling to a common rate, and consistent organization into data structures. Neural recordings include data from approximately 900 worms and 250 uniquely labeled neurons. The connectome data is represented as a graph with attributes for neuron positions, synaptic connections, and additional metadata, facilitating downstream analyses that integrate structure and function.
These datasets are available as open-source resources on the HuggingFace platform, ensuring accessibility for researchers across neuroscience and machine learning. By unifying and releasing these datasets, we aim to catalyze the development of computational models that capture the structure-function relationships in small neural systems, enabling robust cross-lab and cross-animal comparisons.
2 Methods
2.1 Code Repository
We provide an open-source code repository with self-contained ‘preprocess’ and ‘data’ modules that implement the entire preprocessing pipeline we describe in following sections. Our preprocessing is combination of utilities (classes, functions, files and links) that facilitate the extraction and processing of a collection of C. elegans neural and connectome data into standardized formats.
The main classes in the repository are NeuralBasePreprocessor and ConnectomeBasePreprocessor. The former manages processing of neural activity data as measured by calcium fluorescence imaging [33]. The latter manages the processing of synaptic connectivity data as measured by anatomical counts from electron microscopy (EM) [35].
The main functions in the repository are pickle_neural_data() and preprocess_connectome(). The pickle_neural_data() function extracts and processes raw calcium fluorescence data from each of our 11 neural activity source datasets, standardizing them into neural activity matrices that are saved as compressed files for efficient storage and retrieval. The preprocess_connectome() function processes raw connectome data from source files into graph tensors, which are formatted for compatibility with graph neural network (GNN) frameworks such as PyTorch Geometric [21]. This preprocessing step facilitates downstream tasks like connectivity-informed modeling of neural activity. Although there are 10 distinct connectome source files, as detailed in Table 5, all are derived from 4 primary publications [35, 3, 36, 14], ensuring comprehensive coverage of C. elegans neural connectivity data.
Therefore, our code infrastructure allows us to process raw neural and connectome data from different source datasets in a consistent and standard way.
Figure 1: Overview of Neural Activity Datasets. (A) Distribution of the number of worms in each experimental dataset. (B) Number of recorded neurons per worm compared to the total neuronal population of C. elegans. (C) Total duration of recorded neural activity for each dataset. (D) Average duration of recorded activity per worm, with one hour of calcium imaging as a benchmark. (E) Number of sampled time steps per worm, and (F) Sampling intervals for recorded neural activity. The horizontal dashed lines in (E) and (F) indicate the resampled sequence length ($L=314$) and time step ($\Delta t=0.333$ seconds) used in our preprocessing pipeline.
2.2 Calcium Fluorescence Data
We processed neural activity datasets from 11 open-source studies of C. elegans, each measuring calcium fluorescence ($\Delta F/F_{0}$) in subsets of the worm’s neurons under various conditions [14, 5, 18, 37, 9, 15, 12, 10, 11, 13]. These datasets include a variety of experimental protocols, such as freely moving, immobilized, and optogenetically stimulated animals. The number of worms and identified neurons varies across datasets as shown in Table 1. All worms were hermaphrodites at developmental stages no earlier than L4 [29], with the most being in adulthood (Figure 5).
Table 1: Calcium Fluorescence Neural Activity Datasets Metadata. Metadata for calcium fluorescence neural activity datasets collected from various sources. The table includes the number of worms, the average number of labeled and recorded neurons, the range of labeled neurons observed, and summary experimental protocol details.
Index Dataset Name Num. Worms Mean Num. Neurons (labeled, recorded) Num. Labeled Neurons (min, max) Experimental Protocol
1 Kato2015 [10] 12 (42, 127) (31, 51) Wild-type; L4 stage; immobilized in a microfluidic device
2 Nichols2017 [12] 44 (34, 108) (23, 43) Wild-type and npr-1 mutants; Lethargus/pre-lethargus stages; immobilized
3 Skora2018 [15] 12 (46, 129) (39, 55) Wild-type; fasting/starvation conditions; immobilized
4 Kaplan2020 [9] 19 (36, 114) (23, 51) Wild-type; L4 stage; freely moving and immobilized
5 Yemini2021 [37] 49 (110, 125) (33, 179) Transgenic strain OH15500; adult stage; immobilized
6 Uzel2022 [18] 6 (50, 138) (46, 58) Well-fed wild-type; adult stage; immobilized
7 Dag2023 [13] 7 (100, 143) (87, 110) Wild-type; adult stage; freely moving with subsequent NeuroPAL immobilization
8 Atanas2023 [5] 42 (88, 136) (64, 115) Transgenic strain NLS-GCaMP7f; adult stage; freely moving
9 Leifer2023 [14] 103 (69, 122) (26, 102) Wild-type; adult stage; immobilized
10 Lin2023 [11] 577 (8, 8) (1, 22) Wild-type; adult stage; immobilized
11 Venkatachalam2024 [32] 22 (187, 187) (185, 189) Wild-type; adult stage; freely moving
2.2.1 Overview of Neural Activity Datasets
We analyzed neural activity datasets collected from 11 independent studies, which varied in terms of the number of worms recorded, the number of neurons labeled, the duration of recordings, and sampling intervals. Figure 1 provides an overview of these datasets, highlighting their variability in experimental conditions.
2.2.2 Neural Data Processing
The entire pipeline is implemented in the NeuralBasePreprocessor class, which handles dataset-specific preprocessing tasks, including file format loading, trace extraction, and metadata creation.
The neural activity data is preprocessed using a custom pipeline (Figure 2) which follows these steps:
1. Data Download and Extraction. The raw data from each dataset is downloaded from its respective source and extracted. The function download_url() downloads the dataset, and extract_zip() decompresses the data into a local directory while preserving the original folder structure.
2. Custom Dataset Implementations. Each source dataset has a custom class that inherits from the NeuralBasePreprocessor parent class, facilitating data extraction for each unique format. The preprocessor classes ensure that each dataset is loaded correctly, followed by the main preprocessing steps: normalization, smoothing, resampling, and masking. Adding a new source dataset to the pipeline is as easy as defining a new child preprocessor class that implements any loading and processing logic that needs to be customized to that particular source; for example, if the new source’s data is stored in a special file type.
Figure 2: Processed Neural Activity Data Structure. Illustration of the structure of our neural activity data as a time-series matrix, where each column represents a neuron, ordered alphabetically by their canonical names in C. elegans. The calcium signals are normalized, (optionally) smoothed, and resampled to a common time interval ($\Delta t$). A binary mask is generated to indicate which neurons were labeled for each worm.
We represent the neural activity data for a worm $k$ by a matrix $\mathbf{X}^{(k)}\in\mathbb{R}^{T_{k}\times D}$, where $T_{k}$ is the number of time points and $D=300$ is the number of neurons in C. elegans. The neural activity at any time point $t$ is denoted as $\mathbf{X}^{(k)}[t]\in\mathbb{R}^{D}$, which represents the activity snapshot across the $D$ neurons.
3. Neuron Trace Normalization. The first preprocessing step is to normalize ($z$-score) the raw neural activity (i.e. data loaded from the source dataset files) across the neuron dimension. For each neuron trace $i$ in worm $k$, we compute:
$\mathbf{X}^{(k)}_{\text{norm}}=\frac{\mathbf{X}^{(k)}_{\text{raw}}-\mathbf{\mu}}{\mathbf{\sigma}}$
where $\boldsymbol{\mu},\boldsymbol{\sigma}\in\mathbb{R}^{D}$ are vectors of the mean and standard deviation, respectively, for each neuron, computed across the temporal dimension.
4. Smoothing. Optionally, the calcium traces can be smoothed using an exponentially weighted moving average (EWMA). For each neuron, we apply causal smoothing as:
$\mathbf{X}^{(k)}_{\text{smooth}}[t]=\alpha\mathbf{X}^{(k)}_{\text{norm}}[t]+(1-\alpha)\mathbf{X}^{(k)}_{\text{norm}}[t-1]$
where $\alpha$ is a smoothing hyperparameter set in preprocess.yaml. For our released neural dataset, smoothing was not applied ($\alpha=1$) as prior preprocessing had already been performed. Calcium fluorescence data inherently reflects low frequency neural dynamics. Smoothing, which acts further low pass over the raw data, is not needed.
5. Resampling. We resample the data to a fixed time interval $\Delta t$, using linear interpolation for datasets with different temporal resolutions. The new time index $t^{\prime}$ is calculated by:
$t^{\prime}=\left\lfloor\frac{t\cdot\Delta t_{\text{original}}}{\Delta t}\right\rfloor$
The resampled traces $\mathbf{X}_{\text{resampled}}^{(k)}$ are computed as:
$\mathbf{X}_{\text{resampled}}^{(k)}[t^{\prime}]=\mathbf{X}^{(k)}[t]+\left(\mathbf{X}^{(k)}[t+1]-\mathbf{X}^{(k)}[t]\right)\cdot(t^{\prime}-t)$
This ensures consistent temporal resolution across datasets. We use $\Delta t\approx 0.333$ seconds as a compromise to retain high- and low-frequency dynamics.
Importantly, the resampling step is the last preprocessing step necessary to create the neural activity data matrix for each worm:
$\mathbf{X}^{(k)}\leftarrow\mathbf{X}_{\text{resampled}}^{(k)}$
6. Masking and Subsetting. We also create a binary mask for each worm, $\mathbf{M}^{(k)}\in\{0,1\}^{D}$, where $D=300$ represents the total number of canonical neurons in C. elegans. Each element $\mathbf{M}^{(k)}_{i}=1$ indicates that neuron $i$ was labeled in worm $k$, while $\mathbf{M}^{(k)}_{i}=0$ indicates that neuron $i$ was not labeled.
In some of the source datasets we processed there were instances of calcium fluorescence measurements from neurons which were not identified with a label. For simplicity, we will not make a distinction in this work between unlabeled neurons and unmeasured neurons and may use the two terms interchangeably.
The mask $\mathbf{M}^{(k)}$ is stored alongside $\mathbf{X}^{(k)}$. We do not apply the mask to modify the neural data matrix (e.g. zeroing out measurements for unlabeled neurons) but instead retain it as a secondary data structure that can be used in downstream modeling and analysis (e.g., selecting only labeled neurons). Therefore, the pair $(\mathbf{X}^{(k)},\mathbf{M}^{(k)})$ constitute the final output of the neural data processing half of our preprocessing pipeline.
In the next section, we describe the second half the preprocessing pipeline, which unifies many measurements of the of C. elegans into a standardized format.
Algorithm 1 Neural Data Preprocessing Pipeline
[…]
for each dataset $\mathcal{D}_{\mathrm{raw,source}}$ in $\mathcal{D}_{\mathrm{raw,all\_sources}}$ do
     for each worm $k$ in $\mathcal{D}_{i}$ do
         Extract calcium fluorescence traces from source files: $\mathbf{X}^{(k)}_{\text{raw}}$
         Normalize traces: $\mathbf{X}^{(k)}_{\text{norm}}\leftarrow\text{normalize}(\mathbf{X}^{(k)}_{\text{raw}})$
         (Optional) Smooth traces: $\mathbf{X}^{(k)}_{\text{smooth}}\leftarrow\text{smooth}(\mathbf{X}^{(k)}_{\text{norm}},\alpha)$
         Resample traces: $\mathbf{X}^{(k)}_{\text{resample}}\leftarrow\text{resample}(\mathbf{X}^{(k)}_{\text{smooth}},\Delta t)$
         Processed neural data: $\mathbf{X}^{(k)}\leftarrow\mathbf{X}^{(k)}_{\text{resample}}$
         Labeled neuron mask: $\mathbf{M}^{(k)}\leftarrow\{\texttt{boolean}(\mathbf{X}^{(k)}_{j}\neq\emptyset)\mid j\in[1,D]\}$
         Store processed neural data and labeled neuron mask: $\mathbf{X}^{(k)}$ and $\mathbf{M}^{(k)}$
     end for
     Save the processed dataset $\mathcal{D}_{\mathrm{processed,source}}:=\{\left(\mathbf{X}^{(k)},\mathbf{M}^{(k)}\right)\}_{k=1}^{N}$
end for
Compile and store all processed datasets $\mathcal{D}_{\mathrm{processed,all\_sources}}$
2.3 Connectome Graph Data
We processed multiple open-source C. elegans connectome datasets, each capturing the neural wiring diagram, including both chemical synapses and electrical synapse (also referred as gap junctions). Table 5 provides metadata for these datasets, which vary in annotation precision, formats, and the extent of the nervous system covered.
2.3.1 Connectome Representation
The connectome is represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where:
1. Nodes ($\mathcal{V}$): Represent individual neurons.
$\mathcal{V}:=\{\nu_{i}\}_{i=1}^{D},\quad D:=|\mathcal{V}|=300.$
Each node $\nu_{i}$ is associated with attributes, such as:
  • Positional Coordinates: Spatial location of the neuron in body-atlas coordinates $(x,y,z)$ [16].
  • Neuron Class and Type: Categories such as sensory, motor, or interneuron, and neuron-specific types (e.g., ADA, IL2, VD).
  • Neurotransmitter: Identity of neurotransmitters released by the neuron (e.g., acetylcholine, dopamine) [27].
  • Feature Vector ($\mathbf{x}_{i}$): Optionally, numerical data such as calcium fluorescence measurements over time.
2. Edges ($\mathcal{E}$): Represent synaptic and electrical connections between neurons.
:=$\mathcal{E}:=\{e_{ij}\}_{k=1}^{E},\quad E:=|\mathcal{E}|.$
Each edge $e_{ij}$ is associated with an attribute vector $\mathbf{a}_{ij}$, which can include:
  • $c$: Chemical synapse weight.
  • $g$: Gap junction weight.
  • $f$: Functional connectivity strength (optional, based on optogenetic perturbations [14]).
The neighbors of a node $\nu_{i}$ are defined as $\mathcal{N}(\nu_{i}):=\{\nu_{j}:e_{ij}\vee e_{ji}\in\mathcal{E}\}$, where $\vee$ indicates "or", ignoring directionality.
Edges are included in $\mathcal{E}$ if any of their attributes $\mathbf{a}_{ij}$ are nonzero. For example, an edge may represent gap junctions ($\mathbf{a}_{ij}=[0,g]$) or chemical synapses ($\mathbf{a}_{ij}=[c,0]$). In cases where functional connectivity is included, the graph may become fully connected ($|\mathcal{E}|\approx D^{2}$) unless significance thresholds are applied. Chemical and electrical synapse weights are based on counts from EM annotations [35, 3, 36] whereas the functional connectivity is based on signal propagation [14].
Figure 3 illustrates this graph-based representation of the nervous system, highlighting node and edge attributes. Table 5 summarizes the notation used in constructing and describing this graph structure.
2.3.2 Data Extraction and Preprocessing
The data extraction and preprocessing pipeline is implemented in the ConnectomeBasePreprocessor class, which standardizes neuron indices, processes edge attributes, and constructs the final graph tensor Data objects. Each source dataset is handled by a customized subclass since data extraction and preprocessing steps vary by dataset format:
  1. 1.
    Tabular Formats: Most datasets provide neuron pairs, connection types, and weights in a tabular structure. For example, the OpenWorm [22] dataset specifies presynaptic and postsynaptic neurons, connection types (chemical or gap junction), and weights. See Table 5 for an example.
  2. 2.
    Adjacency Matrix Formats: Some datasets use $D\times D$ adjacency matrices, with entries representing synapse counts ([3]) or functional connectivity values ([14]).
  3. 3.
    Supplementing Attributes: Some desirable attributes about the connectome (e.g., neuron positions, neurotransmitters, cell types, etc.) may not available from the source datasets. We obtain these from published research (e.g., neuron positions from [16], neurotransmitters and cell types from [27]) to supplement our graph structures.
Custom classes within the preprocessing pipeline handle dataset-specific formats, ensuring consistent graph tensor outputs.
2.3.3 Graph Tensor Format
To facilitate downstream analyses and modeling, we standardized the connectome data into a graph tensor format compatible with graph neural network (GNN) libraries. Each dataset is transformed into a torch_geometric [21] Data object with the following components:
  • edge_index: Adjacency list encoding $\mathcal{E}$.
  • edge_attr: Attribute vectors $\mathbf{a}_{ij}=[c,g,f,\dots]$ for edges in edge_index.
  • x: Node feature matrix $\mathbf{X}\in\mathbb{R}^{D\times L}$, initialized to zeros or populated with temporal slices of neural activity $\mathbf{X}^{k}[t:t+L]$ for some worm $k$. This data structure is the interface that allows integration of the neural activity data with connectivity data.
  • kwargs: Metadata and node attributes including neuron coordinates, class, type, and neurotransmitter.
This tensor format ensures compatibility with downstream ML models and facilitates integration of functional ($\mathbf{X}$) and structural ($\mathcal{E}$) data. Table 5 provides an overview of the standardized attributes.
2.3.4 Consensus Connectome
To address discrepancies in counts and weights of connections between datasets, we propose a consensus connectome that aggregates data across sources using majority voting. For each edge $e_{ij}$, the consensus weights are:
$c_{\text{consensus}}=\max\{c_{1},c_{2},\dots,c_{n}\},\quad g_{\text{consensus}}=\max\{g_{1},g_{2},\dots,g_{n}\},$
where $c_{i},g_{i}$ are the chemical and gap junction weights from $i$-th connectome source.
In order to quantify the uncertainty of this connectome, we also compute a weighted standard deviation $v$ of each edge across sources and include this as an additional attribute $\mathbf{a}=[c,g,...,v]$:
$v_{\text{consensus}}=\lambda\cdot\operatorname{std}\{c_{1},c_{2},\dots,c_{n}\}+(1-\lambda)\cdot\operatorname{std}\{g_{1},g_{2},\dots,g_{n}\}$
where $\lambda$ is a tuneable hyperparameter determining the relative importance of the chemical versus electrical synapse weights. We set $\lambda=0.9$ reflecting our knowledge that annotations of gap junctions are less certain than those of chemical synapses and that gap junctions are more sparse than chemical synapses in the connectome [3].
Having a single consensus connectome could used to model the average C. elegans nervous system, which has a stereotyped neurodevelopmental trajectory [35], or to quantify the inter-animal variability in network weights, which may reflect environmental influences. Figure 5 visualizes the consensus connectome of the 22 amphid chemosensory neurons [11].
Figure 3: Graph-Based Representation of the Connectome. Graph-based structure used for the C. elegans connectome data. Nodes represent neurons, with node attributes (e.g., position, class) and optional features. Directed edges capture connections, with edge attributes (e.g., chemical and gap junction weights).
3 Discussion
In this paper we present a normalized and standardized datasets for C. elegans neural activity and connectomes. We have created and consolidated code blocks which can be used to load neural activity and connectome data from a variety of different source datasets with different file formats and subsequently preprocess this data. The code is composed of class definitions for each source dataset inheriting from a common NeuralBasePreprocessor class and a ConnectomeBasePreprocessor class along with helper functions which create class objects and execute class functions.
The homogenized neural and connectome datasets provide a unique opportunity to explore the relationship between neural dynamics and structural connectivity in a fully mapped nervous system. While our datasets provide a strong foundation for building models that link neural dynamics to structure, the limitations of both the neural activity data and the connectome must be taken into account when using these datasets. There are several important limitations and methodological choices that must be considered when interpreting the data and applying it in both biological and computational models. We address most most of these in the next few paragraphs.
One of the primary limitations of our neural activity dataset arises from the use of calcium fluorescence imaging as the method for measuring neural activity. Calcium imaging, while valuable for capturing broad patterns of activity across large populations of neurons, is an indirect measure of neural activity and introduces a low-pass filtering effect [6]. Calcium signals lag behind the actual electrical activity due to the slower dynamics of calcium ion concentration changes compared to the fast voltage changes associated with action potentials. This delay can be particularly problematic for studies investigating fast synaptic interactions, as important rapid neural dynamics may be lost. Thus, while calcium imaging is useful for measuring slow, global brain states, it cannot fully capture the complexity of rapid information processing that may be critical in certain neural circuits [33, 20]. This highlights the need to complement calcium-based measurements with other techniques, such as those utilizing voltage-sensitive indicators, to gain a more comprehensive understanding of neural activity. However, calcium dynamics may be sufficient for the study of neural signaling in C. elegans as most of its neurons are nearly isopotential in the steady state [23].
While the connectome of C. elegans provides a complete structural blueprint of the nervous system, there are uncertainties and limitations inherent in structural connectomics. One major limitation is the variability in synaptic strengths between neurons, both across individuals and across the lifespan of a single individual [36]. The structural connectome is essentially a static snapshot, and we do not yet fully understand how synaptic strengths might fluctuate under different physiological conditions [14]. Moreover, the connectome data used in this study does not come from the same animals used for neural activity measurements, creating a mismatch between the dynamics and connectivity measurements. This creates both an opportunity and a limitation. The opportunity arises from the fact that any robust relationship between structure and function to be discovered from this data must generalize across different age and phenotype-matched animals. The limitation is that this may preclude the development of individual-specific models which are desirable for understanding learning and memory, for example. While having the C. elegans connectome [35, 3] provides a comprehensive structural map of synaptic connections, understanding the relationship between connectivity and dynamics still requires careful interpretation. Additionally, there is evidence for wireless connections between neurons [14] due to neuromodulation by way of neuropeptides which can bind to nearby and distant neuron GPCR sites triggering secondary messaging pathways [1, 31].
We opted not to apply smoothing to the neural activity traces for two principal reasons. First, it is likely that the original data creators have already undertaken a degree of preprocessing on their raw microscopy or imaging data, which often includes smoothing techniques to enhance signal quality. Introducing additional smoothing could risk over-processing the data, thereby obscuring critical neural dynamics. Second, the calcium fluorescence signals utilized in our analysis are inherently low-pass filtered representations of the underlying neural activity, a characteristic of the imaging modality itself. This filtering effectively diminishes high-frequency components, which are vital for capturing rapid neural events. By refraining from further smoothing, we aim to preserve the integrity and richness of the original data, allowing for a more nuanced understanding of the biological phenomena under investigation.
Resampling was chosen as a strategy to make the data comparable across experiments with different temporal resolutions. We resampled the neural activity traces to a common time step of approximately $\Delta t\approx 0.333$ seconds, which was chosen as a balance between temporal resolution and minimizing data loss due to downsampling from experiments with higher frame rates. However, this decision introduces its own limitations. There was a tradeoff between attenuating high-frequency information from datasets with higher sampling rates versus incorrectly hallucinating data from datasets with lower sampling frequencies. For experiments that originally had much higher temporal resolution, downsampling risks losing potentially valuable high-frequency dynamics. Conversely, for lower-resolution datasets, resampling can artificially inflate the temporal resolution without actually adding more meaningful information [11]. This was a necessary compromise to allow for cross-study comparisons but should be considered when interpreting the results.
One important clarification about our work needs to made regarding our use of $300$ for the number of neurons in C. elegans, whereas readers may be more familiar the number $302$ from prior literature. We drop the bilateral pair of canal-associated neurons (CAN) from our datasets because they completely lack chemical synapses and are now widely classified as ‘end-organs’ rather than neurons [17, 3]. The absence of synaptic polarity information, as mentioned earlier, limits the ability of models to fully capture the functional consequences of the structural connections. Moreover, the used of calcium measurements might bias neural activity models towards capturing slow, large-scale patterns at the expense of finer, faster neural interactions. Researchers should be mindful of these constraints when developing models and consider integrating additional multimodal datasets to build more comprehensive models [8].
Despite these limitations, the C. elegans neural dataset represents a valuable resource for computational research, particularly in developing biologically grounded models of neural dynamics. For instance, graph neural networks (GNNs) [34] are well-suited to handle the structural connectivity data, while time-series models like structured state space models [25] can be applied to the neural activity data to capture temporal dependencies. These models could inform the development of foundation models that generalize to larger and more complex systems. Moreover, the datasets’ open availability facilitates collaborative efforts in both neuroscience and AI, encouraging the development of new techniques and models that can integrate structure and function in novel ways.
In conclusion, while our homogenized datasets offer significant opportunities for advancing our understanding of small neural systems, its limitations should be carefully considered in both biological and computational contexts. Calcium imaging, noisy connectomics, and the standardization choices we made each come with trade-offs that impact how the data can be used. As neural imaging technologies continue to improve and new datasets are integrated, we expect that this dataset will serve as a foundation for building more complete models of neural function, bridging the gap between simple nervous systems like C. elegans and the more complex neural architectures found in larger organisms.
4 Conclusion
We present a standardization protocol implemented in a code repository for the preprocessing and integration of 11 C. elegans calcium fluorescence datasets and 3 primary connectome sources [35, 3, 36]. The code repository contains preprocessing classes that load and preprocess the different source datasets, handling both calcium fluorescence and connectome data. Each dataset has a class implementation that inherits from a respective parent class, which has been adapted to accommodate varying raw data formats. Once loaded, the calcium fluorescence data is further masked, normalized, resampled, and smoothed.
This unified dataset offers an accessible and standardized resource, streamlining access to data recorded under diverse experimental conditions and formats. The datasets are available for public access on Hugging Face: (1) neural data, and (2) connectome data. Additionally, the full code repository containing the preprocessing code can be accessed on GitHub.
Such standardization facilitates the development of models for the entire C. elegans nervous system, allowing researchers to study the interplay between structure, via the connectome, and function, via calcium fluorescence data. Through this integration, we hope to provide further insights into how neural function emerges from structural connectivity and how these dynamics might generalize to more complex nervous systems.
The open-source nature of this resource aims to further contribute to both neuroscience and AI research, accelerating efforts to model not only C. elegans but also organisms with larger and more complex nervous systems. By releasing this dataset and accompanying code, we hope to advance the growing initiative toward neural structure-function modeling and the discovery of principles of neural computation across species.